OpenStudy (anonymous):

Solve for v 9v^2=30v-25

OpenStudy (anonymous):

Can some one explain how to do these? if you subtract 9v^2 from left to right i get \[0=-9v^2+30v-25\] and i don't know what's next.

OpenStudy (anonymous):

Solve the quadratic. Either by factoring or using the formula.

OpenStudy (anonymous):

polpak i don't know how to do either one of them... by the way which formula?

OpenStudy (anonymous):

It doesn't look like a nice one to factor offhand so I suggest using the formula. But if you haven't learned it you may have to use the hard way.

OpenStudy (anonymous):

\[ax^2 + bx + c =0 \implies x = {-b \pm \sqrt{b^2-4ac} \over 2a}\]

OpenStudy (anonymous):

quadratic formula?

OpenStudy (anonymous):

Lets see if it's easier to factor than I think.

OpenStudy (anonymous):

Oh, ! think it is!

OpenStudy (anonymous):

-25? i got 5/3 when i solved it the formula

OpenStudy (anonymous):

Yeah, it's a perfect square.

OpenStudy (anonymous):

If we multiply your equation by -1 it makes more sense.

OpenStudy (anonymous):

a and c are 3. b and d are -5

OpenStudy (anonymous):

so the leading coefficient would be 9 not -9 i see.

OpenStudy (anonymous):

We have (3v - 5)(3v - 5)

OpenStudy (anonymous):

Foiling that out we should get what we started with.

OpenStudy (anonymous):

If you don't know the formula though you have to figure out how to factor by hand.

OpenStudy (anonymous):

i know the formula but you can't take a square root of a negative number. this is what i mean. \[x=\frac{5}{3}\pm \sqrt{-1800}\]

OpenStudy (anonymous):

oops. it's \[\frac{ \sqrt{-1800}}{18}\]

OpenStudy (anonymous):

\[\sqrt{b^2 - 4(a)(c)} = \sqrt{30^2 - 4(-9)(-25)}\]

OpenStudy (anonymous):

So it's \[\sqrt{900 - 900} = 0\]

OpenStudy (anonymous):

yeah but weren't we supposed to multiply the whole equation by -1? so we get \[\sqrt{-30^2-4(9)(25)}\]

OpenStudy (anonymous):

That was for factoring.

OpenStudy (anonymous):

And even if we had it wouldn't matter because \[(-30)^2 = 30^2 = 900\]

OpenStudy (anonymous):

i got what you got also. it ends up only \[\frac{5}{3}\]

OpenStudy (anonymous):

Yeah, you will get the same answer even if you multiply by -1 because the b is squared and the a times the c will keep the same sign.

OpenStudy (anonymous):

\[\sqrt{(-30)^2-4(9)(25)} = \sqrt{900 - 900} = 0\]

OpenStudy (anonymous):

ok a*c=-9 ad+cb=30 b*d=-25 you wrote this above how do you get your answer or how do you factor out a polynomial with the formula that you've also wrote \[(av^2+b)(cv^2+d)\]

OpenStudy (anonymous):

Sorry my formula should have been: \[(av+b)(cv+d) = 9v^2 - 30v + 25\] Therefore \[a\times c = 9\]\[b \times d = 25\]\[(a\times c) + (b\times d) = -30\]

OpenStudy (anonymous):

But there's not enough equations to solve it properly. If you know for sure that a, b, c, and d are integers you can make some good guesses though.

OpenStudy (anonymous):

For example, b and d are either 25 and 1, or 5 and 5. c and a are either 3 and 3 or 9 and 1. And for either pair the signs for that pair have to be the same.

OpenStudy (anonymous):

so that way is guess and check practically? i see

OpenStudy (anonymous):

Ack

OpenStudy (anonymous):

That last equation is wrong.

OpenStudy (anonymous):

Should be \[(a\times d) + (c\times b) = -30\]

OpenStudy (anonymous):

Those coefficients come from the foiling process.

OpenStudy (anonymous):

So we can see that if ad = cb we would have 2ad = -30 so ad = -15 so 5 and 3 are good choices.

OpenStudy (anonymous):

So one of the pairs is negative, the other is positive and we're good.

OpenStudy (anonymous):

ok thanks for your time polpak. got to run be back in 20 min. maybe some more question will arouse.

OpenStudy (anonymous):

sure thing =)